OpenStudy (anonymous):

the coordinates of parallelogram UVWZ are U(a,0), W(c-a,b), and Z(c,0). find the coordinates of V without using any new variables. Can someone help me out please and thank you

OpenStudy (anonymous):

the diagonals of paralleogram bisects each other so mid point of UW and VZ are same. use this condition to find V

OpenStudy (anonymous):

mid point of UW=(c/2,a/2) so mid point of VZ=(c/2,a/2) let v=(x,y) so mid point of VZ=((c+x)/2,y/2) now (c+x)/2=c/2 x=0 y/2=a/2 y=a V=(0,a)

OpenStudy (anonymous):

thanks for the help

OpenStudy (dumbcow):

one thing i noticed, the midpoint should be (c/2, b/2) yielding V= (0,b) Also you could use the fact that the distance is same for UZ and VW distance = c-a since they are parallel the y coordinate remains b (c-a) - (c-a) = 0 thus V= (0,b)